Follow the 5 steps below to complete this problem.
Step 1:
Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following:
Drive a boat or Drive a jet ski
Step 2:
Select a current speed of the water in mph.
Step 3:
Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2.
Step 4:
Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current).
Step 5:
Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

Answer by ankor@dixie-net.com(22740)

Follow the 5 steps below to complete this problem.
Step 1:
Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following:
Drive a boat or Drive a jet ski
Step 2:
Select a current speed of the water in mph.
Step 3:
Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2.
Step 4:
Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current).
:
Mortimer pedals his paddle boat up the stream to visit Minnie the Moocher in 2 hrs
After a disgusting visit, he heads down stream, and was glad to be be home in 1 hr.
the speed of the current was 2 mph
:
Let s = his pedal speed in still water
:
write a dist equation; dist = time * speed
:
Upstream dist = downstream dist
2(s-2) = 1(s+2)
2s - 4 = s + 2
2s -s = 2 + 4
s = 6 mph is his speed in still water
:
:
Confirm the solution by finding the distance, they should be equal
2(6-2) = 8 mi
1(6+2) = 8 mi